OpenStudy (anonymous):

How many H atoms are present in a sample of 3.5 grams of C6H12O6?

TheSmartOne (thesmartone):

How about trying the Chemistry section..

OpenStudy (anonymous):

Still technically a math question, but yes, probably better suited for Chemistry. At any rate... \[\begin{array}{c|c} \text{Element}&\text{Molecular mass (amu)}\\ \hline C&\sim12\\ H&\sim1\\ O&\sim16 \end{array}\] The molar mass of \(C_6H_{12}O_6\) is \[(6\times12+12\times1+6\times16)\approx180\frac{\text{g}}{\text{mol}}\] In 3.5 g of \(C_6H_{12}O_6\), you have \[3.5\text{ g}\times\frac{1}{180}\frac{\text{mol}}{\text{g}}\approx0.019\text{ mol}\] Of the 0.019 mole of \(C_6H_{12}O_6\), hydrogen takes up half, so there is about 0.0097 mole of \(H\) in the sample. Multiplying by Avogadro's number will give you the number of hydrogen atoms.